OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Add 9 to both sides of the equation Add 5 to both sides of the equation Subtract 5 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (anonymous):

Anyone please

OpenStudy (anonymous):

@jim_thompson5910 @jdoe0001

OpenStudy (anonymous):

@StudyGurl14 @superhelp101

OpenStudy (anonymous):

@cookiemonster2412

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@jdoe0001 please help

OpenStudy (jdoe0001):

\(\bf x^2 + 18x + 76 = 0\implies (x^2 + 18x) + 76 = 0 \\ \quad \\ (x^2 + 18x+{\color{red}{ \square ?}}^2) + 76 = 0\) so... any ideas?

OpenStudy (jdoe0001):

recall http://openstudy.com/study#/updates/5450237de4b0c26d0cb2a454 the middle term value

OpenStudy (anonymous):

Dont know, if you figure out the answer I can tell you how to do it

OpenStudy (anonymous):

Because I am totally confused

OpenStudy (jdoe0001):

well... you may want to review your perfect square trinomial section then

OpenStudy (anonymous):

Ok, but could you show me the answer?

OpenStudy (anonymous):

I promise I will review it

OpenStudy (jdoe0001):

well... that nullifies the idea behind the exercise

OpenStudy (anonymous):

Ok I will answer your questions

OpenStudy (anonymous):

I reviewed it too

OpenStudy (anonymous):

@jdoe0001 you there?

OpenStudy (anonymous):

Oh i think it is 9

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Bro

OpenStudy (anonymous):

@jdoe0001 ? You there?

OpenStudy (anonymous):

@jdoe0001 please help me

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@PinkCam88 @purplepiggy171 @radar@robtobey @jdoe0001 @Walkerman15 @blahhedup101

OpenStudy (anonymous):

please someone help

OpenStudy (anonymous):

Anyone who helps will get a medal

OpenStudy (litening1337):

IDK.

OpenStudy (neonumbrella5115):

Well, I never learned this formula. So, sorry I can't help.

OpenStudy (neonumbrella5115):

@HelpBlahBlahBlah @doulikepiecauseidont @cody7744 @Walkerman15

OpenStudy (neonumbrella5115):

@Bobtoktok Sorry I couldn't help, but I hope those people can! ;)

OpenStudy (anonymous):

I'm sorry I don't know